OpenStudy (anonymous):

Solve the equation -50= p - 43

OpenStudy (anonymous):

add 43 to both sides: p=-7

OpenStudy (anonymous):

if you are happy with the answer please click on good answer to close the question

OpenStudy (asadkarim7):

-50+43= p-43+43 p=-7